# Bilateral facial reduction: qualification-free subdifferential calculus and exact duality in convex analysis

### v1

@@ Abstract
This paper introduces a geometric framework to extend convex analysis results beyond standard qualification conditions such as intersecting relative interiors of domains. We define the joint facial subspace $T$ as the span of the face of CDC-D generated by 0, with its affine translation $T_{a}$ containing $C\cap D$. The intersections $C\cap T_{a}$ and $D\cap T_{a}$ are faces of the original sets, establishing $T_{a}$ as a bilateral facial reduction with parallels to Borwein-Wolkowicz facial reduction in conic programming. When reducing the domains of two convex functions to their joint facial subspace $T_{a}$, their relative interiors then always intersect, enabling unqualified application of classical theorems via localization. Key generalizations include subdifferential additivity, normal cones of intersections, a subdifferential chain rule, attained infimal convolution for $(f+g)^{*},$ and an exact Fenchel-Rockafellar dual. We characterize $T$ via generated faces and computationally through an iterative process converging in at most $n$ steps in $\mathbb{R}^{n}$. Proofs are self-contained and introduce novel concepts like facial subspaces and nested normals to describe convex boundaries and the lattice of faces.
@@ 1 Introduction
In convex analysis, the effective domain of the sum of two convex functions is the intersection of their effective domains. Many central results, including the summation rule of subdifferentials, the convex conjugate of sums of functions, or the strong duality of the Fenchel-Rockafellar dual, exclude the case where the domains of the convex functions intersect only on their relative boundaries. The conditions excluding such cases are called qualification conditions, or constraint qualifications depending on the context. One such condition, widespread in the seminal work [14] by Rockafellar, is that for two functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
$\operatorname{relint}(\operatorname{dom}(f))\cap\operatorname{relint}(\operatorname{dom}(g))\neq\emptyset.$
For instance, this condition appears in [14, Theorem 23.8], which establishes the summation rule of subdifferentials:
$\partial(f+g)(x)=\partial f(x)+\partial g(x).$
A second example is the strong duality of the Fenchel-Rockafellar dual: for $A\in\mathbb{R}^{m\times n}$, denoting by $f^{*}$ the convex conjugate of $f$,
$\inf_{x\in\mathbb{R}^{n}}f(x)+g(Ax)=-\inf_{y\in\mathbb{R}^{m}}[f^{*}(A^{T}y)+g^{*}(-y)].$
In [14, Corollary 31.2.1], this is guaranteed to hold if
$\operatorname{relint}(\operatorname{dom}(f))\cap A^{-1}(\operatorname{relint}\operatorname{dom}(g))\neq\emptyset,$
where $A^{-1}(\cdot)$ is the preimage for the linear map induced by the matrix $A$. This requirement is often called a constraint qualification, and is strikingly similar to the qualification condition for the subdifferential sum rule.
These and a number of other results share similar conditions. The object of this work is to generalize such results beyond the regime where these conditions hold. Specifically, in Section 5 we find generalizations for the two aforementioned results, as well as a few others: the attainment of infimal convolution as the convex conjugate of the sum of two convex functions, the normal cone of the intersection of two convex sets, and a chain rule for the subdifferential. There are many reasons such generalizations are worthwhile, and we list two here. From a theoretical perspective, extending such results makes the theory more elegant and intuitive. A second motivation is that such generalizations could lead to improved methods in optimization, as did facial reduction in conical programming (see [3]).
Qualification conditions are not merely theoretical artifacts; they are supported by counter examples, which indicates that these conditions reflect more or less accurately the true scope of the statements. We consider one such counter-example for the sum rule of subdifferentials. Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise. This function has effective domain $\mathbb{R}_{+}$. Adding it to the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, the sum rule states that $\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset$, when in fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$. The relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect. The qualification condition does not hold, so there is no contradiction with [14, Theorem 23.8], yet this still shows that the sum rule does not in fact apply to the sum of any two convex functions. Thus, an accurate generalization of the sum rule must offer a different formula for the subdifferential of the sum of two convex functions, and such a formula should ideally reduce to the classical sum rule when qualification conditions are satisfied. We provide such a formula in Theorem 5.1.
Our approach to relaxing qualification conditions is to focus on the geometry of the effective domain of the convex functions. Qualification conditions fail when relative interiors do not intersect, in which case only the boundaries of the domains meet. We show that the boundaries of two convex sets always meet in a structured way: each set presents a convex face to the other, and both of these faces are revealed by the intersection with a common affine subspace, which we name the joint facial subspace $T_{a}$. The main property of the joint facial subspace is that given convex sets $C,D\subseteq\mathbb{R}^{n}$, with $T_{a}$ defined as in Definition 4.1, whenever $C\cap D\neq\emptyset$,
@@
The results which shows this is Corollary 4.3. Therefore, qualification conditions always hold when localizing to the joint facial subspace. Roughly, we generalize results in convex analysis by replacing convex functions $f$ by their localized version $f+\iota_{T_{a}}$. The modified results then hold even without qualification conditions. And when qualification conditions do hold for the original results, then $T_{a}=\mathbb{R}^{n}$ (see Proposition 5.1), and the additional indicator function can be ignored.
For instance, in Theorem 5.1 we state that
$\partial(f+g)(x)=\partial(f+\iota_{T_{a}})(x)+\partial(g+\iota_{T_{a}})(x),$
without any condition on $f$ and $g$ other than they be convex functions. This modification works for the previously encountered counter example. In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint facial subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
@@
which is indeed $\partial(f+g)(x)$. One way to interpret our result is that there are subgradients that are “hiding” in the faces of the domains of the convex functions. Intersecting with the joint facial subspace reveals these hidden gradients.
The central theoretical concept introduced in this work is the joint facial subspace. We therefore focus much of our efforts in describing this object. This is the focus of Section 4, where we define the joint facial subspace in Definition 4.1 , provide two characterizations and state core property, that qualification conditions hold when localizing to it. Given two convex sets $C,D\subseteq\mathbb{R}^{n}$, the joint facial subspace is defined from the Minkowski difference $C-D=\{c-d\mid c\in C,d\in D\}$. In the first characterization, Theorem 4.1, it is made clear how the joint facial subspace relates to $C$ and $D$ individually. The third characterization is computational in nature, and describes how the joint facial subspace can be found from a number of steps that is bounded by the ambient dimension.
Another important contribution of the present work is in our proof techniques. In Section 6, we build a theoretical framework with nested normals (Definition 6.3) and facial subspaces (Definition 6.2), two objects which together describe the lattice of faces of a convex set. Their combination forms a dual object which is more descriptive than the standard normal cone (see Theorem 6.2). Our theory is therefore at least in part a theory about the boundary of convex sets. From this toolkit, we are able to describe the faces of differences of convex sets CDC-D in Section 6.2, from which we derive the various properties of the joint facial subspace in Section 4. These objects will be further explored and refined in future work.
The structure of the paper goes as follows. In Section 2 surveys prior works and Section 3 introduces some notation. Section 4 is as described above, with proofs deferred to Section 7. In Section 5, we provide generalizations of central results in convex analysis, together with their proofs, which use Corollary 4.3 from Section 4. In Section 6, we give a motivated introduction to our mathematical framework, introducing key concepts. In Section 7, we prove the results in Section 4 by using tools from Section 6.
@@ 2 Prior works
The existing research which is the most related to this work is Facial Reduction (FR) in conic programming  [2, 1, 3], which helped guide the development of the theory in this work. We briefly describe this method. Given a feasible cone and affine constraints, Slater’s condition (that the interior of the feasible cone has at least one point satisfying the constraints), FR narrows the ambient space of the problem to the span of a “minimal” face of the feasible cone. After narrowing the problem via this “preprocessing” step, the problem is guaranteed to satisfy Slater’s condition.
Our joint facial subspace does something similar, but for any two intersecting convex sets. We find a “minimal” face of the difference of the sets, CDC-D (see Definition 4.1), and this yields the joint facial subspace. The computation of the joint facial subspace can be similarly viewed as a processing step, after which the problem is reduced to the joint facial subspace. As specified in [3], FR can be applied with finite-step algorithms, with a number of steps bounded by the ambient dimension. This is very similar to our second characterization of the joint facial subspace, Theorem 7.1.
A result which is related (where the relation is examined in [11]), is the Extended Dual System of Ramana, a dual program for semidefinite programming for which strong duality always holds [13]. The corresponding result in our work is the exact Fenchel-Rockafellar dual in Theorem 5.5.
From these strong parallels, we conjecture that many aspects of Facial Reduction and the Extended Dual System can be recovered—and potentially generalized—from results in this work.
Prior efforts to obtain a qualification-free sum rule for subdifferentials have developed powerful limiting approaches that achieve exact characterizations without traditional qualifications. For instance, $\epsilon$-subdifferentials define approximate subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential as $\varepsilon$ approaches zero [4, 12, 10]. Similarly, fuzzy sum rules approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [6, 5]. These methods, leveraging local neighborhood behavior, are related to ideas in variational analysis, and offer versatile tools for theoretical and computational analysis across general spaces. They can often be implemented with fixed tolerances for practical computation, though the exactness relies on infinite limiting processes to eliminate approximation errors. Though our approach is tailored to pairs of functions, our formula for the subdifferential of a sum of convex functions is exact, and can be computed in finite steps.

### v3

# Qualification-free convex analysis via the joint supporting subspace

@@ Abstract
In convex analysis, qualification conditions (also termed constraint qualifications) help avoid pathological behavior at domain boundaries. In this work we remove the need for such conditions by localizing to an affine subspace—the joint supporting subspace—that contains the feasible region and ensures qualification conditions hold after localization.
Our theory generalizes Borwein and Wolkowicz’s facial reduction beyond conic programs to convex programs of the form $f(x)+g(Ax)$. Intuitively, the joint supporting subspace corresponds to a bilateral facial reduction between any two convex sets. It enables simple qualification-free generalizations for a host of central results of convex analysis. These include: an exact Fenchel-Rockafellar dual; Karush-Kuhn-Tucker (KKT) optimality conditions; attained infimal convolution for convex conjugates; subdifferential sum and chain rules; and a characterization of the normal cones of the intersection of two convex sets. All generalizations reduce seamlessly to their original formulations when qualification conditions hold.
We offer a number of characterizations for the joint supporting subspace, one of which is constructive. Our proofs are self-contained, and introduce a novel theoretical framework consisting of nested normals and supporting subspaces, which simultaneously describe both the boundary of convex sets and the lattice of faces.
@@ 1 Introduction
In convex analysis, many results depend on qualification conditions, of which one prominent example found in the seminal work [rockafellarConvexAnalysis1970] by Rockafellar, is that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$,
$\operatorname{relint}\operatorname{dom}(f)\cap\operatorname{relint}\operatorname{dom}(g)\neq\emptyset,$ (1)
where $\operatorname{relint}$ denotes the relative interior. This condition ensures that the feasible region $\operatorname{dom}(f+g)$ is not contained in the relative boundary of either $\operatorname{dom}(f)$ or $\operatorname{dom}(g)$. When Equation 1 does not hold, the geometry of the boundary of the convex sets $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$ becomes important, and requires special treatment. Such cases are typically excluded from consideration with qualification conditions because of the added complexity, and also because such cases are often avoidable in practice.
The sum rule of subdifferentials,
$\partial(f+g)(x)=\partial f(x)+\partial g(x),$
is guaranteed by [rockafellarConvexAnalysis1970, Theorem 23.8], subject to Equation 1 (amongst other possible sufficient conditions). Other results of convex analysis with similar conditions include, e.g., the subdifferential chain rule, attainment of the infimal convolution in convex conjugates, strong duality of the Fenchel-Rockafellar dual, and Karush-Kuhn-Tucker (KKT) optimality conditions. For results concerned with optimization problems, this type of condition is usually called constraint qualifications.
Qualification conditions are not merely theoretical artifacts; they are justified by counterexamples, and therefore reflect more or less accurately the true scope of applicability of the associated rules. We introduce one such counterexample for the sum rule of subdifferentials, to later show how our theory resolves this case. Take the function $f(x)=-\sqrt{x}$ for $x\geq 0$ and $f(x)=\infty$ otherwise. Consider the indicator function $g(x)=\iota_{\mathbb{R}_{-}}(x)$, which is 0 on $\mathbb{R}_{-}$ and $\infty$ elsewhere. Naive application of the sum rule yields
$\partial(f+g)(0)=\partial f(0)+\partial g(0)=\emptyset+\mathbb{R}_{+}=\emptyset.$
In fact, $\partial(f+g)(0)=\partial\iota_{0}(0)=\mathbb{R}$ so the sum rule does not hold. There is no conflict with [rockafellarConvexAnalysis1970, Theorem 23.8] because qualification conditions are violated; the relative interiors of the domains are $(-\infty,0)$ and $(0,\infty)$, and do not intersect. The standard sum rule fails in this case, though we will show that it is recovered after localization to the joint supporting subspace.
We tackle the problem of qualification conditions directly by developing geometric results describing the boundary of convex sets, allowing us to generalize a collection of results from convex analysis in one fell swoop in Section 5, with the classical formulations trivially recovered when qualifications hold (see Proposition 5.1).
The key geometric insight is as follows. For any two intersecting convex sets $C,D\subseteq\mathbb{R}^{n}$ in $\mathbb{R}^{n}$, there exists a special affine subspace $T_{a}$—the joint supporting subspace—that contains $C\cap D$ and that is uniquely determined by $C$ and $D$, i.e., $T_{a}:=T_{a}(C,D)$. We provide multiple characterizations of the joint supporting subspace in Section 4. Supporting subspaces (see Definition 6.1) are a generalization of supporting hyperplanes, first defined, to our knowledge, in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4]. As intersections with supporting hyperplanes reveal exposed faces of convex sets, so do intersections with supporting subspaces reveal faces—which need not be exposed. As the name suggests, $T_{a}$ is a supporting subspace of both $C$ and $D$, which means that two faces of $C$ and $D$ are simultaneously revealed by intersection with $T_{a}$. Each of the convex sets presents a face to the other, and they meet within $T_{a}$. The utility of $T_{a}$ in convex analysis stems from two simple facts:
  1. 1.
    $C\cap D\subseteq T_{a}.$
  2. 2.
@@
We prove these facts in Corollary 4.1 and Corollary 4.5 respectively. These two properties can be understood as follows: when reducing the ambient space to $T_{a}$, the the reduced ( localized, intersected) sets $C^{\prime}:=C\cap T_{a}$ and $D^{\prime}:=D\cap T_{a}$ preserve the original intersection ($C^{\prime}\cap D^{\prime}=C\cap D$) while satisfying the qualification condition $\operatorname{relint}(C^{\prime})\cap\operatorname{relint}(D^{\prime})\neq\emptyset$. Similarly, for convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, with $T_{a}:=T_{a}(\operatorname{dom}(f),\operatorname{dom}(g))$, the localized functions $f^{\prime}:=f+\iota_{T_{a}},g^{\prime}:=g+\iota_{T_{a}}$ preserve the original sum ($f^{\prime}+g^{\prime}=f+g$) while satisfying Equation 1. Standard convex analysis results can then be applied to the localized functions $f^{\prime}$ and $g^{\prime}$, while still accurately describing the original sum $f+g$. For instance, in Corollary 5.1 we state that
$\partial(f+g)(x)=\partial f^{\prime}(x)+\partial g^{\prime}(x),$ (2)
without any condition on $f$ and $g$ other than they be proper convex functions. A host of other generalizations are obtained via minor variations on this procedure in Section 5.
This modification to the sum rule resolves the previously encountered counterexample. In that problem, where $f(x)=-\sqrt{x}$ and $g(x)=\iota_{(-\infty,0]}$, the joint supporting subspace from Definition 4.1 is $T_{a}=\{0\}$, and so
@@
which indeed matches $\partial(f+g)(x)$.
1.1 Facial reduction and computational outlook
As we show in Section 5.3.1, our theory is a generalization of facial reduction, a preprocessing technique with strong theoretical guarantees in conic programming. Facial reduction reduces the problem to the smallest face of the conic constraint that contains the feasible set [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981]. After reducing to this face, Slater’s condition is provably satisfied. As mentioned in [drusvyatskiyManyFacesDegeneracy2017], facial reduction leads to both a reduction in the size of the problem and improved numerical stability, and has consequently been implemented in solvers for semi-definite programs (e.g., in the MOSEK solver [permenterSolvingConicOptimization2017, fribergFacialReductionMOSEK]). The impact of facial reduction in computation inspires hope that our theory could lead to improvements in optimization algorithms in non-conic settings where constraint qualifications may fail.
Where facial reduction reduces to the minimal face of a cone, we reduce jointly to the smallest faces of two convex sets, both selected by intersection with the single joint supporting subspace. For simplicity, we frame our results as a reduction of the relevant objects (convex sets or convex functions), while keeping the ambient space $\mathbb{R}^{n}$ fixed. Our results are however naturally suited to a reduction of the ambient space to the joint supporting subspace, i.e., the joint supporting subspace would then become the new ambient space (see, e.g., Theorem 5.1). This alternative mathematical formulation would better model a reduction of the problem in a preprocessing step. Reduction of the ambient space to the joint supporting subspace in a preprocessing step may seem counterintuitive, as results in convex analysis are commonly specified only for real vector spaces, not affine spaces. This discrepancy can be mitigated by translating the affine subspace jointly with all the objects it contains until it reaches origin. But even this minor adjustment seems redundant, because, we believe, affine spaces seem to be the natural home for many results in convex analysis.
In full abstraction, an affine space is a pair $(L,U)$ consisting of a set $L$ and a real vector space $U$ (the tangent space), equipped with an addition operation $L\times U\to L$ with the property that $\forall l_{1},l_{2}\in L$ there exists a unique $u\in U$ such that $l_{1}+u=l_{2}$. For $L$ an affine subspace of $\mathbb{R}^{n}$, the tangent space is $U=L-L=\{l_{1}-l_{2}\mid l_{1},l_{2}\in L\}$. Results in convex analysis can then be generalized by distinguishing tangent vectors from points in the affine space. For example, that affine subspaces are closed under convex combinations follows from reformulating $\lambda x+(1-\lambda)y$ for $x,y\in L$ as $y+\lambda(x-y)$, by noticing that $y\in L$ and that $\lambda(x-y)$ is in the tangent space LLL-L. The tangent space appears in our definition for the separation of convex sets (Definition 4.3), and justifies our repeated use of self-differences of sets SSS-S (see, e.g., Definition 6.2, Corollary 4.1, or Definition 4.2).
1.2 Theoretical contribution
An important contribution of the present work is the theoretical framework that we build in Section 6. The central objects are nested normals (Definition 6.3) and supporting subspaces (Definition 6.2), which describe local subsets of the lattice of faces of a convex set. Their combination forms a mathematical object that is an exact dual for convex cones (see Theorem 6.2). As such, the nested normals can also be understood as a local description of the boundary of convex sets, an idea that we discuss in Section 6. With these objects, we derive results describing the intersection of convex sets, as well as the facial structure of differences of convex sets. Notable technical results include Theorem 6.3, Theorem 6.4, and Theorem 6.6. We therefore contribute a geometric framework describing the boundaries, intersections, and differences of convex sets.
1.3 Paper outline
Section 2 surveys prior works, notably facial reduction in conic programming, limiting approaches to exact subdifferential calculus, and the lexicographic characterization of the faces of convex sets. In Section 3 we introduce some notation. In Section 4 we state our results describing the joint supporting subspace, with proofs deferred to Section 7. There, we define the joint supporting subspace in Definition 4.1, characterize it with a distinct analytical formula in Theorem 4.1, and with a constructive iterative procedure in Corollary 5.7 which terminates in a number of steps no more than the ambient dimension. In Theorem 4.2 we state the main property of the joint supporting subspace: that qualification conditions hold when localizing to it. In Section 5, we state and prove generalizations of central results in convex analysis. In Section 5.3.1, we show that results of facial reduction in conic programming can be recovered from our results. In Section 6, we provide a narrative, motivated introduction to our mathematical framework, introducing key concepts and proving associated results along the way. In Section 7, we prove the results of Section 4 with the tools from Section 6.
@@ 2 Prior works
Constraint qualifications (CQs) in finite dimensions evolved from Slater’s condition [slaterLagrangeMultipliersRevisited1959], later weakened to the algebraic Mangasarian-Fromovitz CQ [mangasarianFritzJohnNecessary1967] and to the geometric Abadie CQ [abadieNonlinearProgramming1967]. The weak Slater condition, introduced by Rockafellar [rockafellarConvexAnalysis1970], uses the notion of relative interiors. Subsequent relaxed CQs include the constant rank constraint qualification condition [janinDirectionalDerivativeMarginal1984], further relaxed in [qiConstantPositiveLinear2000, andreaniRelaxedConstantPositive2012], and a CQ by [robinsonFirstOrderConditions1976], which takes the form of a stability condition now known as linear metric subregularity. We note that in his seminal work [rockafellarConvexAnalysis1970], Rockafellar used the same qualification conditions for optimization problems and subdifferential calculus, which enabled our unified treatment of optimization and subdifferential calculus.
There has been two major research directions generalizing results beyond qualification conditions. The geometric stream, to which our work belongs, focuses on the geometry of constraints and feasible sets. The sequential stream considers limits of increasingly accurate approximations, with exact results in the limit, and is related to variational analysis [mordukhovich_VariationalAnalysisApplications_2018, rockafellarVariationalAnalysis1997]. The geometric stream focuses on constraint qualifications for optimization problems, while the sequential stream emphasizes subdifferential calculus, with extensions to sequential optimality conditions. Our work extends the geometric stream to subdifferential calculus. We survey qualification-free results in finite dimensions that are comparable to our results in Section 5.
2.1 Geometric stream
In the setting of inequality-constrained convex programs, the Fritz-John condition [johnExtremumProblemsInequalities2014] is an optimality condition that holds without qualification conditions. Unlike the KKT condition, it is not a sufficient condition for optimality, and is satisfied either when KKT conditions hold, or when some degeneracy in the constraints is detected, akin to applying the KKT conditions with a suitable constraint qualification.
A qualification-free necessary and sufficient condition was formulated by [ben-talCharacterizationOptimalityConvex1976], and extended to the nonsmooth case by [wolkowiczGeometryOptimalityConditions1980]. These works add a term to the KKT condition to deal with pathological cases; see [dharaOptimalityConditionsConvex2011, Chapter 6] for a treatment. Our qualification-free KKT condition in Corollary 5.6 is formulated for programs of the form $f(x)+g(Ax)$, instead of inequality-constrained convex programs, and uses faces of convex sets instead of an additional term.
Facial Reduction (FR) generalized preprocessing ideas from linear programming to the setting of conic programming [massamOptimalityConditionsConeConvex1979, borweinFacialReductionConeconvex1981, borweinRegularizingAbstractConvex1981], with subsequent developments providing efficient computational methods for the reduction. An early method iteratively reduces the dimension of the problem [borweinRegularizingAbstractConvex1981, luoDualityResultsConic1997, wakiFacialReductionAlgorithms2013] until the minimal face is reached. Ideas of facial reduction can alternatively be framed with exact dual problems [borweinCharacterizationOptimalityAbstract1981], with important applications to semidefinite programming [ramanaExactDualityTheory1997, patakiStrongDualityConic2013]. For in-depth literature reviews of facial reduction and its numerous applications, we refer the reader to [tuncelStrongDualityMinimal2012, drusvyatskiyManyFacesDegeneracy2017, permenterReductionMethodsSemidefinite2017].
Many results in our work have analogues in facial reduction. For instance, the notion of a minimal face appears in different forms in Definition 4.1 and Theorem 4.1. The iterative reduction procedure of [borweinRegularizingAbstractConvex1981] finds its analogue in Theorem 7.1. We derive an exact Fenchel-Rockafellar dual in Corollary 5.5, echoing exact duals in conic programming. Finally, our qualification-free optimality condition in Corollary 5.6 generalizes similar results in the conic setting.
2.2 Sequential stream
The sequential approach to relaxing qualification conditions began with subdifferential calculus. One approach is that of $\varepsilon$-subdifferentials, which defines subgradients satisfying the subgradient inequality up to an $\varepsilon$-error, recovering the exact subdifferential in the limit of vanishing $\varepsilon$ [hiriarturruty_SubdifferentialCalculusUsing_1993, penotSubdifferentialCalculusQualification1996, mordukhovich_VariationalAnalysisApplications_2018]. A second, similar approach is that of fuzzy sum rules, which approximate the subdifferential by perturbing the point $x$ at which the subdifferential is taken, deriving exact additivity rules in the limit of vanishing perturbations [jules_FormulasSubdifferentialsSums_2002, ioffeFuzzyPrinciplesCharacterization1998].
These ideas were then applied to optimality conditions, where the condition of optimality takes the form of the existence of some sequence converging to the optimal point. One approach is to consider sequences of points converging to the point of interest [thibaultSequentialConvexSubdifferential1997], similarly to fuzzy sum rules, while another is analogous to $\varepsilon$-subdifferentials [jeyakumarInequalitySystemsGlobal1996, jeyakumarNewSequentialLagrange2003] relating them to the epigraphs of conjugate functions.
2.3 Theory of convex faces